# High-inclination Centaur reservoirs beyond Neptune

### v1

@@ Abstract
Context. Numerical simulations of the 4.5 Gyr past evolution of high-inclination Centaurs showed they originated from orbits beyond Neptune that were perpendicular to the Solar System’s invariable plane in a region called the polar corridor. The existence of the polar corridor is explained by the Tisserand inclination pathways followed by Neptune-crossing objects in the three-body problem. Recently, a study of Centaur injection in the three-body problem showed that Neptune-crossing transneptunian objects (TNOs) in the polar corridor with semi-major axes in the range [40:160] au have dynamical times that exceed the Solar System’s age suggesting the possible presence of long-lived reservoirs that produce high-inclination Centaurs.
Aims. We aim to demonstrate numerically the existence of such reservoirs in the Solar System by simulating the TNOs’ time-forward evolution under the gravitational perturbations of the giant planets, the Galactic tide and passing stars. We also aim to assess the efficiency of Centaur injection as a function of the initial inclination and determine if high-inclination Centaurs may be produced by low inclination reservoirs.
Methods. The motion of the giant planets, TNOs and passing stars is simulated using the IAS15 $N$-body numerical integrator of the REBOUND package. The Galactic tide is included using the REBOUNDx package. Two TNO orbit types are considered in the semi-major axis range [40:140] au: cold TNOs with circular orbits and hot TNOs with perihelion range [32:50] au. The TNOs Tisserand parameters $T$ with respect to Neptune are taken in the range $2\leq T\leq 2.8$ corresponding to inclinations far from Neptune in the range $8^{\circ}\leq I_{\infty}\leq 135^{\circ}$ in order to examine Centaur injection at low and high initial inclinations.
Results. We find that TNO reservoirs in the semi-major axis range [50:140] au are long-lived and their populations peak at $T=0.5$ and T=1.5T=-1.5. Saturn is found to induce secondary structures in the polar corridor by holding the perihelia of a fraction of high-inclination reservoir material. We find that the Centaur inclination at minimum semi-major axis depends linearly on the Tisserand parameter regardless of the initial semi-major axis. Its amplitude shows that low inclination reservoirs such as the early protoplanetary disk are unlikely to produce high-inclination Centaurs in contrast to reservoirs in the polar corridor.
Conclusions. We identified the likely location of the closest reservoirs to Neptune populated by TNOs captured in the early Solar System that produce high-inclination Centaurs. The Legacy Survey of Space and Time of the Vera Rubin Observatory will be able to constrain the reservoirs’ extent and population size.
@@ 1 Introduction
Solar System Centaurs with inclinations larger than $60^{\circ}$ are conventionally known as high-inclination Centaurs. Like low inclination Centaurs, they originate from planet-crossing orbits beyond Neptune but unlike low inclination Centaurs, their numbers cannot be explained by the protoplanetary disk in the early Solar System (Fernandez16; Fraser22; KaibVolk22). High-resolution simulations of the past orbits of 19 real high-inclination Centaurs over 4.5 Gyr showed that they tend to be polar with respect to the Solar System’s invariable plane accumulating in the so-called polar corridor that extends beyond Neptune’s orbit (NamouniMorais18b; NamouniMorais20b) (hereafter Papers I and II). The polar corridor’s existence is rooted in the conservation of the Tisserand parameter of Neptune-crossing TNOs that ensures that their dynamical evolution follows Tisserand inclination pathways over Gyr-timescales regardless of whether time flows forward or backward (Namouni22) (hereafter Paper III).
@@
The semi-major axis spread of 4.5-4.5 Gyr high-inclination Centaur clone orbits in the polar corridor extends from the scattered disk’s inner edge to the Oort cloud indicating a wide region of origin (Papers I and II). The Oort cloud, in particular, is a known possibility (Brasser12; Ito24). The presence of Oort cloud matter at 4.5-4.5 Gyr requires it to have been captured by the Solar System from the material in the stellar cluster of its birth (Fernandez00; Levison10; Brasser12; Jilkova16; Hands19; Kaib19). Studies of the relaxation of the early flat protoplanetary disk through the orbital instability phase of the Solar System’s planets demonstrated that the Oort cloud formed through the scattering of disk matter later than 4.5-4.5 Gyr cannot explain the numbers of high-inclination Centaurs (Kaib19; Nesvorny19). Orbits of high-inclination Centaurs may be generated by hypothetical planets in the outer Solar System from an extended Kuiper belt (Batygin16b; Lykawka23) but these models neglect the stellar environment of the Solar System known to induce instability over Gyr timescales (Kaib25).
@@
In this paper, we examine a new possible radial origin suggested by the semi-major axis spread of 4.5-4.5 Gyr high-inclination Centaur clone orbits in the polar corridor. The 4.5-4.5 Gyr semi-major axis probability distribution functions were found to peak outside Neptune’s orbit in the semi-major axis range [70:100] au (Paper II). These peaks may indicate a stable region where TNO reservoirs are currently present in the polar corridor and have been supplying high-inclination Centaurs to the giant planets’ domain. As a first step in understanding the dynamical origin of the semi-major axis distribution peaks, we studied the dynamics of Neptune-crossing TNOs in the Sun-Neptune-TNO three-body problem (Namouni24) (hereafter Paper IV). This first step is important as the dynamics of Neptune-crossing TNOs is determined to a great extent by the conservation of the Tisserand parameter that gives rise to the concept of Tisserand inclination pathways. A Tisserand inclination pathway describes how orbital inclination evolves as Neptune is approached, and how the TNO is possibly injected as a Centaur before it eventually recedes from the planet to the outer Solar System if it does not experience collision or ejection. The Tisserand inclination pathway depends only on the TNO’s Tisserand parameter and the planet’s semi-major axis. In particular, it is independent of time. This makes the Tisserand parameter the perfect tool to probe original inclinations. The study of the TNO injection process in Paper IV, as a function of the Tisserand parameter, revealed the existence of a stable region in the polar corridor with semi-major axis range [40:160] au where the dynamical time of planet-crossing TNOs exceeds the Solar System’s age suggesting the possible presence of long-lived Centaur-supplying TNO reservoirs. The fact that the stable region found in the three-body problem using time-forward integration is consistent with the 4.5-4.5 Gyr semi-major axis distribution peaks of Paper II obtained using time-backward integration in the presence of the four giant planets is encouraging. If the existence of this stable region is confirmed numerically in a comprehensive time-forward simulation of the Solar System, it will indicate the likely location of the closest reservoirs to Neptune populated by TNOs that were captured in the polar corridor 4.5 Gyr ago and that produce high-inclination Centaurs. In this paper, we provide such confirmation.
@@
In Section 2, we describe the initial conditions used in this work and their relationship to the stable region found in Paper IV. In particular, parameter space is enlarged beyond the three-body stable region’s Tisserand parameter range and beyond the perihelion range of the past orbits of high-inclination Centaurs of Papers I and II. In Section 3, we describe the simulation method used to follow the time-forward evolution of TNO reservoirs subject to the perturbations of the four giant planets, the Galactic tide and passing stars. The reservoirs’ state at the simulation’s end is discussed as a function of the initial semi-major axis and Tisserand parameter in Section 4. The reservoirs’ stability is confirmed as well as Neptune’s role in providing Tisserand inclination pathways for the reservoirs’ material to follow. It is also found that Saturn produces secondary structures in the polar corridor as it retains a fraction of TNOs perihelia near its orbit. In Section 5, Centaur injection is described. In particular, the number of injected Centaurs is found to depend linearly on the Tisserand parameter. In Section 6, the features of Centaurs orbits are described. Additional linear relations with respect to the Tisserand parameter are found, for the Centaur minimum semi-major axis, for the Centaur perihelion at minimum semi-major axis and the Centaur inclination at minimum semi-major axis. The latter explains why high-inclination Centaurs are unlikely to be produced from low inclination reservoirs such as the early protoplanetary disk. Section 7 contains concluding remarks. In the remainder of this paper, The Tisserand parameter is specified with respect to Neptune unless stated otherwise.
@@ 2 Reservoirs’ initial extent
Figure 1: Initial state of the TNO reservoirs for different Tisserand parameters $T$ indicated in the top right corner of each panel. Hot TNOs are shown in red and cold TNOs are shown in blue. The solid orange and blue curves in the inclination panels are respectively the Tisserand inclination pathway (2) and the initial inclination of circular orbits (5). In the eccentricity panels, the curves indicate perihelia that match the four giant planets’ semi-major axes.
We build the reservoirs’ initial state using three results from the 4.5 Gyr time-backward numerical simulation of 19 real high-inclination Centaurs using $\sim 20$ million clones (Papers I and II). The first result is the Centaurs’ past polar state. Starting with initial inclinations extending over $62^{\circ}$$173^{\circ}$], the 4.5-4.5 Gyr end state inclinations extended over $60^{\circ}$$90^{\circ}$] with a standard deviation $\sim 10^{\circ}$ beyond Neptune’s orbit. An analytical analysis of the three-body problem (Paper III) demonstrated that this clustering is the manifestation of the Tisserand parameter’s conservation that in turn makes Centaurs and TNOs follow the Tisserand inclination pathway:
@@
$\mbox{where}\ \ \ \ T$ $=$ $\frac{a_{p}}{a_{0}}+2\left[\frac{a_{0}(1-e_{0}^{2})}{a_{p}}\right]^{\frac{1}{2}}\cos I_{0},$ (2)
$e_{0}$ and $I_{0}$ are the asteroid’s initial eccentricity and inclination and $a_{0}$ ad $a_{p}$ are the asteroid’s and the planet’s semimajor axes. The Tisserand inclination pathway is independent of time and perihelion and defines the polar corridor as the extent in aaII-space between $I_{\rm Tiss}(a,2,a_{p})$ and $I_{\rm Tiss}(a,-1,a_{p})$. This corresponds approximately to the inclination range of the 4.5-4.5 Gyr orbits of high-inclination Centaurs.
The second result is the extent of the 4.5-4.5 Gyr semi-major axis distributions peaks in the region [70:100] au. This extent is smaller than the three-body stable region (Paper IV), [40:160] au, where TNO dynamics was followed at Neptune’s collision singularity. The discrepancy in the inner boundary has two reasons. First, the giant planets’ perturbations are stronger at the inner edge of the three-body stable region. Secular perturbations from the giant planets, in particular, increase instability as they may raise significantly the eccentricity of planet-crossing and non-crossing TNOs. Second, the initial conditions at Neptune’s singularity include, for a given Tisserand parameter, TNOs with small perihelia that would have crossed the unmodeled orbits of Jupiter, Saturn and Uranus and became unstable. In contrast, Papers I and II indicated that past high-inclination Centaur orbits in the polar corridor did not have arbitrary perihelia but they clustered mainly in the vicinity of Neptune’s orbit. The discrepancy in the outer boundary is unclear. Despite the proximity of the 100 au boundary to the location of the outer 1:6 mean motion resonance with Neptune, it was shown in Paper IV that resonances do not fully explain the stability of high-inclination Neptune-crossing TNOs.
The Tisserand parameter range of the three-body stable region where the dynamical time exceeds 4.5 Gyr is [1-1:0] corresponding to an inclination far from Neptune $90^{\circ}\leq I_{\infty}\leq 110^{\circ}$ where $I_{\infty}$ is given as (Paper III):
@@
In Papers III and IV, we demonstrated that a TNO may be injected inside Neptune’s orbit only if its Tisserand parameter $T\geq-1$. Once injected, its Centaur inclination becomes retrograde if $T\leq 2$ and remains prograde otherwise. The T=1T=-1 limit is specific to the three-body problem and stems from the Tisserand inclination pathway. The pathway guides the TNO towards the planet up to a reflection semi-major axis that depends on $T$ and Neptune’s semi-major axis $a_{\rm N}$ given as:
@@
The TNO then reverses its radial motion and moves away from the planet. The limit T=1T=-1 is when reflection occurs at Neptune’s semi-major axis. For T<1T<-1, the TNO is reflected outside the planet’s orbit and may not become a Centaur. The addition of Jupiter, Saturn and Neptune to the simulation modifies the three-body injection process and a larger $T$-range is necessary. The Tisserand parameter of a TNO that may cross a planet’s orbit must satisfy $\sqrt{8}\leq T\leq\sqrt{8}\sim 2.82$ (Paper III). To probe the transfer of prograde orbits to high-inclination retrograde orbits, the reservoirs’ Tisserand parameter range is enlarged to [2-2:$2.5$] with a 0.5-step from the three-body problem stability range of [1[-1:0]. The additional value $T=2.8$ ($I_{\infty}\sim 8^{\circ}$) was included to provide a means of comparison with low-inclination dynamics.
We examine two orbit types in the semi-major axis range [40:140] au with a 10 au-step: hot and cold TNOs. Hot TNOs are test particles with eccentric orbits and perihelia in the range [32:50] au with a 2 au-step [^1] For $a=40$ au, 32 $\leq q\leq 38 au$. These reservoirs could be part of the material captured by the Solar System in its birth cluster. The perihelion range is based on the third result of the time-backward simulations that showed polar corridor orbits’ perihelia clustered at Neptune’s orbit. Enlarging the perihelion range beyond Neptune’s semi-major axis explores whether the giant planets’ combined perturbations produce high-inclination Centaurs from an initial population of non-planet-crossing TNOs.
@@
For each set ($a$,$T$,$q$), 100 orbits are generated by drawing the orbital angles, mean longitude, longitude of ascending node and argument of perihelion, from a uniform distribution. As the Tisserand inclination pathway is independent of any angles, the 100 orbits may evolve on the same pathway in the three-body problem. Furthermore, regardless of the initial $a$ and $q$, ensembles with the same $T$ value may evolve on the same Tisserand inclination pathway.
Figure 1 shows the reservoirs’ initial extent as a function of eccentricity, inclination and semi-major axis for T=2,1.5,1, 0, 0.5, 2, 2.5T=-2,\ -1.5,\ -1,\ 0,\ 0.5,\ 2,\ 2.5 and $2.8$. Hot and cold TNOs are shown in red and blue respectively. Larger perihelia imply larger (smaller) inclinations than the Tisserand pathway’s for (retrograde) prograde motion. This deviation from the inclination pathway increases with the initial semi-major axis and has a limit value, $I_{\rm circ}$, for circular orbits given as:
@@
and shown in the inclination panels of Figure 1 as the blue curve along with the Tisserand inclination pathway (2) as the orange curve.

### v2

@@ Abstract
Context. Numerical simulations of the 4.5 Gyr past evolution of high-inclination Centaurs show that they originated from orbits beyond Neptune that were perpendicular to the Solar System’s invariable plane in a region called the polar corridor. The existence of the polar corridor is explained by the Tisserand inclination pathways followed by Neptune-crossing objects in the three-body problem. Recently, a study of Centaur injection in the three-body problem has shown that Neptune-crossing trans-Neptunian objects (TNOs) in the polar corridor with semimajor axes in the range [40:160] au have dynamical times that exceed the Solar System’s age, suggesting the possible presence of long-lived reservoirs that produce high-inclination Centaurs.
Aims. We aim to numerically demonstrate the existence of such reservoirs in the Solar System by simulating the TNOs’ time-forward evolution under the gravitational perturbations of the giant planets, the Galactic tide, and passing stars. We also aim to assess the efficiency of Centaur injection as a function of the initial inclination and determine if high-inclination Centaurs can be produced by low-inclination reservoirs.
Methods. The motion of the giant planets, TNOs, and passing stars was simulated using the IAS15 $N$-body numerical integrator of the REBOUND package. The Galactic tide was included using the REBOUNDx package. Two TNO orbit types were considered in the semimajor axis range [40:140] au: cold TNOs with circular orbits and hot TNOs with a perihelion range of [32:50] au. The TNO Tisserand parameters, $T$, with respect to Neptune were taken in the range [2-2:2.8] which corresponds to inclinations far from Neptune in the range $8^{\circ}$$135^{\circ}$], in order to examine Centaur injection at low and high initial inclinations.
Results. We find that TNO reservoirs in the semimajor axis range [50:140] au are long-lived and their populations peak at $T=0.5$ and T=1.5T=-1.5. Saturn is found to induce secondary structures in the polar corridor by holding the perihelia of a fraction of high-inclination reservoir material. We find that the Centaur inclination at minimum semimajor axis depends linearly on the Tisserand parameter regardless of the initial semimajor axis. Its amplitude shows that low-inclination reservoirs such as the early protoplanetary disk are unlikely to produce high-inclination Centaurs, in contrast to reservoirs in the polar corridor.
Conclusions. We have identified the likely location of the closest reservoirs to Neptune populated by TNOs captured in the early Solar System that produce high-inclination Centaurs. The Legacy Survey of Space and Time of the Vera Rubin Observatory will be able to constrain the reservoirs’ extent and population size.
@@ 1 Introduction
Solar System Centaurs with inclinations greater than $60^{\circ}$ are conventionally known as high-inclination Centaurs. Like low-inclination Centaurs, they originate from planet-crossing orbits beyond Neptune, but unlike low-inclination Centaurs, their numbers cannot be explained by the protoplanetary disk in the early Solar System (Fernandez16; Fraser22; KaibVolk22). High-resolution simulations of the past orbits of 19 real high-inclination Centaurs over 4.5 Gyr have shown that they tend to be polar with respect to the Solar System’s invariable plane, accumulating in the so-called polar corridor that extends beyond Neptune’s orbit (NamouniMorais18b; NamouniMorais20b, hereafter Papers I and II). The polar corridor’s existence is rooted in the conservation of the Tisserand parameter of Neptune-crossing trans-Neptunian objects (TNOs), which ensures that their dynamical evolution follows Tisserand inclination pathways over gigayear timescales regardless of whether time flows forward or backward (Namouni22, hereafter Paper III).
@@
The semimajor axis spread of 4.5-4.5 Gyr high-inclination Centaur clone orbits in the polar corridor extends from the scattered disk’s inner edge to the Oort cloud, indicating a wide region of origin (Papers I and II). The Oort cloud, in particular, is a known possible source (Brasser12; Ito24). The presence of Oort cloud matter at 4.5-4.5 Gyr requires it to have been captured by the Solar System from the material in the stellar cluster of its birth (Fernandez00; Levison10; Brasser12; Jilkova16; Hands19; Kaib19). Studies of the relaxation of the early flat protoplanetary disk through the orbital instability phase of the Solar System’s planets have demonstrated that the Oort cloud component that was formed through the scattering of disk matter later than 4.5-4.5 Gyr cannot explain the numbers of high-inclination Centaurs (Kaib19; Nesvorny19). Orbits of high-inclination Centaurs may be generated by hypothetical planets in the outer Solar System from an extended Kuiper belt (Batygin16b; Lykawka23), but such models neglect the stellar environment of the Solar System known to induce instability over gigayear timescales (Kaib25).
@@
In this paper we examine a new possible radial origin suggested by the semimajor axis spread of 4.5-4.5 Gyr high-inclination Centaur clone orbits in the polar corridor. The 4.5-4.5 Gyr semimajor axis probability distribution functions were found to peak outside Neptune’s orbit in the semimajor axis range [70:100] au (Paper II). These peaks potentially indicate a stable region where TNO reservoirs are currently present in the polar corridor and have been supplying high-inclination Centaurs to the giant planets’ domain. As a first step in understanding the dynamical origin of the semimajor axis distribution peaks, we studied the dynamics of Neptune-crossing TNOs in the Sun-Neptune-TNO three-body problem (Namouni24, hereafter Paper IV). This first step is important as the dynamics of Neptune-crossing TNOs is determined to a great extent by the conservation of the Tisserand parameter that gives rise to the concept of Tisserand inclination pathways. A Tisserand inclination pathway describes how orbital inclination evolves as Neptune is approached, and how the TNO is possibly injected as a Centaur before it eventually recedes from the planet to the outer Solar System if it does not experience collision or ejection. The Tisserand inclination pathway depends only on the TNO’s Tisserand parameter and the planet’s semimajor axis. In particular, it is independent of time. This makes the Tisserand parameter the perfect tool for probing original inclinations. The study of the TNO injection process in Paper IV, as a function of the Tisserand parameter, revealed the existence of a stable region in the polar corridor with a semimajor axis range of [40:160] au where the dynamical time of planet-crossing TNOs exceeds the Solar System’s age, suggesting the possible presence of long-lived Centaur-supplying TNO reservoirs. The fact that the stable region found in the three-body problem using time-forward integration is consistent with the 4.5-4.5 Gyr semimajor axis distribution peaks obtained in Paper II using time-backward integration in the presence of the four giant planets is encouraging. In this study we numerically confirm the existence of this stable region in a comprehensive time-forward simulation of the Solar System; in doing so, we determined the likely location of the closest reservoirs to Neptune populated by TNOs that were captured in the polar corridor 4.5 Gyr ago and that produce high-inclination Centaurs.
@@
In Sect. 2 we describe the initial conditions used in this work and their relationship to the stable region found in Paper IV. In particular, the parameter space is enlarged beyond the three-body stable region’s Tisserand parameter range and beyond the perihelion range of the past orbits of high-inclination Centaurs studied in Papers I and II. In Sect. 3 we describe the simulation method used to follow the time-forward evolution of TNO reservoirs subject to the perturbations of the four giant planets, the Galactic tide, and passing stars. The reservoirs’ state at the simulation’s end is discussed as a function of the initial semimajor axis and Tisserand parameter in Sect. 4. The reservoirs’ stability is confirmed, as is Neptune’s role in providing Tisserand inclination pathways for the reservoirs’ material to follow. It is also found that Saturn produces secondary structures in the polar corridor as it retains a fraction of TNO perihelia near its orbit. In Sect. 5 Centaur injection is described. In particular, the number of injected Centaurs is found to depend linearly on the Tisserand parameter. In Sect. 6 the features of Centaurs orbits are described. Additional linear relations with respect to the Tisserand parameter are found: for the Centaur minimum semimajor axis, for the Centaur perihelion at minimum semimajor axis, and for the Centaur inclination at minimum semimajor axis. The third explains why high-inclination Centaurs are unlikely to be produced from low-inclination reservoirs such as the early protoplanetary disk. Section 7 contains concluding remarks. In the remainder of this paper, the Tisserand parameter is specified with respect to Neptune unless stated otherwise.
@@ 2 Reservoirs’ initial extent
Figure 1: Initial state of the TNO reservoirs for different Tisserand parameters, $T$, as indicated in the top-right corner of each panel. Hot TNOs are shown in red and cold TNOs in blue. The solid orange and blue curves in the inclination panels are, respectively, the Tisserand inclination pathway (2) and the initial inclination of circular orbits (5). In the eccentricity panels, the curves indicate perihelia that match the four giant planets’ semimajor axes.
We built the reservoirs’ initial state using three results from the 4.5 Gyr time-backward numerical simulation of 19 real high-inclination Centaurs using $\sim 20$ million clones (Papers I and II). The first result is the Centaurs’ past polar state. Starting with initial inclinations extending over $62^{\circ}$$173^{\circ}$], the 4.5-4.5 Gyr end state inclinations extended over $60^{\circ}$$90^{\circ}$] with a standard deviation $\sim 10^{\circ}$ beyond Neptune’s orbit. An analytical analysis of the three-body problem (Paper III) demonstrated that this clustering is the manifestation of the Tisserand parameter’s conservation, which makes Centaurs and TNOs follow the Tisserand inclination pathway:
@@
$\mbox{where}\ \ \ \ T$ $=$ $\frac{a_{p}}{a_{0}}+2\left[\frac{a_{0}(1-e_{0}^{2})}{a_{p}}\right]^{\frac{1}{2}}\cos I_{0}.$ (2)
Here $e_{0}$ and $I_{0}$ are the asteroid’s initial eccentricity and inclination and $a_{0}$ ad $a_{p}$ are the asteroid’s and the planet’s semimajor axes. The Tisserand inclination pathway is independent of time and perihelion and defines the polar corridor as the extent in aaII-space between $I_{\rm Tiss}(a,2,a_{p})$ and $I_{\rm Tiss}(a,-1,a_{p})$. This corresponds approximately to the inclination range of the 4.5-4.5 Gyr orbits of high-inclination Centaurs.
The second result is the extent of the 4.5-4.5 Gyr semimajor axis distributions peaks in the region [70:100] au. This extent is smaller than the three-body stable region (Paper IV), [40:160] au, where TNO dynamics was followed at Neptune’s collision singularity. The discrepancy in the inner boundary has two reasons. First, the giant planets’ perturbations are stronger at the inner edge of the three-body stable region. Secular perturbations from the giant planets, in particular, increase instability as they can raise significantly the eccentricity of planet-crossing and non-crossing TNOs. Second, the initial conditions at Neptune’s singularity include, for a given Tisserand parameter, TNOs with small perihelia that would have crossed the un-modeled orbits of Jupiter, Saturn and Uranus and became unstable. In contrast, Papers I and II indicated that past high-inclination Centaur orbits in the polar corridor did not have arbitrary perihelia but they clustered mainly in the vicinity of Neptune’s orbit. The discrepancy in the outer boundary is unclear. Despite the proximity of the 100 au boundary to the location of the outer 1:6 mean motion resonance with Neptune, it was shown in Paper IV that resonances do not fully explain the stability of high-inclination Neptune-crossing TNOs.
The Tisserand parameter range of the three-body stable region where the dynamical time exceeds 4.5 Gyr is [1-1:0] corresponding to an inclination far from Neptune $90^{\circ}\leq I_{\infty}\leq 110^{\circ}$ where $I_{\infty}$ is given as (Paper III)
@@
In Papers III and IV, we demonstrate that a TNO can be injected inside Neptune’s orbit only if its Tisserand parameter $T\geq-1$. Once injected, its Centaur inclination becomes retrograde if $T\leq 2$ and remains prograde otherwise. The T=1T=-1 limit is specific to the three-body problem and stems from the Tisserand inclination pathway. The pathway guides the TNO toward the planet up to a reflection semimajor axis that depends on $T$ and Neptune’s semimajor axis $a_{\rm N}$ given as
@@
The TNO then reverses its radial motion and moves away from the planet. The limit T=1T=-1 is when reflection occurs at Neptune’s semimajor axis. For T<1T<-1, the TNO is reflected outside the planet’s orbit and cannot become a Centaur. The addition of Jupiter, Saturn and Neptune to the simulation modifies the three-body injection process and a larger $T$-range is necessary. The Tisserand parameter of a TNO that can cross a planet’s orbit must satisfy $\sqrt{8}\leq T\leq\sqrt{8}\sim 2.82$ (Paper III). To probe the transfer of prograde orbits to high-inclination retrograde orbits, the reservoirs’ Tisserand parameter range is enlarged to [2-2:$2.5$] with a 0.5-step from the three-body problem stability range of [1[-1:0]. The additional value $T=2.8$ ($I_{\infty}\sim 8^{\circ}$) was included to provide a means of comparison with low-inclination dynamics.
We examined two orbit types in the semimajor axis range [40:140] au with a 10 au-step: hot and cold TNOs. Hot TNOs are test particles with eccentric orbits and perihelia in the range [32:50] au with a 2 au-step [^1] For $a=40$ au, 32 $\leq q\leq 38 au$. These reservoirs could be part of the material captured by the Solar System in its birth cluster. The perihelion range is based on the third result of the time-backward simulations that showed polar corridor orbits’ perihelia clustered at Neptune’s orbit. Enlarging the perihelion range beyond Neptune’s semimajor axis explores whether the giant planets’ combined perturbations produce high-inclination Centaurs from an initial population of non-planet-crossing TNOs.
@@
For each set ($a$,$T$,$q$), 100 orbits are generated by drawing the orbital angles, mean longitude, longitude of ascending node and argument of perihelion, from a uniform distribution. As the Tisserand inclination pathway is independent of any angles, the 100 orbits evolve on the same pathway in the three-body problem. Furthermore, regardless of the initial $a$ and $q$, ensembles with the same $T$ value evolve on the same Tisserand inclination pathway.
Figure 1 shows the reservoirs’ initial extent as a function of eccentricity, inclination and semimajor axis for T=2,1.5,1, 0, 0.5, 2, 2.5,T=-2,\ -1.5,\ -1,\ 0,\ 0.5,\ 2,\ 2.5, and $2.8$. Hot and cold TNOs are shown in red and blue, respectively. Larger perihelia imply larger (smaller) inclinations than the Tisserand pathway’s for (retrograde) prograde motion. This deviation from the inclination pathway increases with the initial semimajor axis and has a limit value, $I_{\rm circ}$, for circular orbits given as
@@
and shown in the inclination panels of Fig. 1 as the blue curve along with the Tisserand inclination pathway (2) as the orange curve.